You find a correlation between enthusiasm for Cookies and happiness of r = .8. For 12 participants, your sum of cookie enthusiasm (i.e., sum of X) was 120, (sum of X-squared = 1,308), and your sum for happiness (sum of Y) was 144 (sum of Y-squared = 1,920).


Based on this information, calculate the regression equation that predicts happiness as a function of cookie enthusiasm.

Make predictions about happiness values for cookie enthusiasm scores of 3 and 10.
Show all relevant work for all components of this problem.
Answer #1

Solution:

Given:

x = cookie enthusiasm

y = happiness

Correlation coefficient = r = 0.8

n = 12

1 = 120

\sum x^{2} = 1308

\sum y = 144

\sum y^{2} = 1920

Calculate the regression equation that predicts happiness as a function of cookie enthusiasm.

y =b_{0}+b_{1}\times x

where

b_{1} = r \times \frac{s_{y}}{s_{x}}

s_{y}=\sqrt{\frac{\sum y^{2}-(\sum y)^{2}/n}{n-1}}

s_{x}=\sqrt{\frac{\sum x^{2}-(\sum x)^{2}/n}{n-1}}

thus

s_{y}=\sqrt{\frac{\sum y^{2}-(\sum y)^{2}/n}{n-1}}

s_{y}=\sqrt{\frac{1920-(144)^{2}/12}{12-1}}

s_{y}=\sqrt{\frac{1920-1728 }{11}}

s_{y}=\sqrt{\frac{192 }{11}}

s_{y}=\sqrt{ 17.454545 }

s_{y}= 4.177864

and

s_{x}=\sqrt{\frac{\sum x^{2}-(\sum x)^{2}/n}{n-1}}

s_{x}=\sqrt{\frac{1308-(120)^{2}/12}{12-1}}

s_{x}=\sqrt{\frac{1308-1200 }{11}}

s_{x}=\sqrt{\frac{108 }{11}}

s_{x}=\sqrt{ 9.818182 }

s_{x}=3.133398

thus

b_{1} = r \times \frac{s_{y}}{s_{x}}

b_{1} = 0.8 \times \frac{4.177864 }{3.133398 }

b_{1} = 1.066667

and

b_{0}=\bar{y}-b_{1}\times \bar{x}

where

\bar{y}=\frac{\sum y}{n}=\frac{144}{12} = 12

\bar{x}=\frac{\sum x}{n}=\frac{120}{12}=10

thus

b_{0}=\bar{y}-b_{1}\times \bar{x}

b_{0}=12- 1.066667 \times 10

b_{0}=12- 10.66667

b_{0}=1.33333

thus the regression equation that predicts happiness as a function of cookie enthusiasm is:

y =b_{0}+b_{1}\times x

\mathbf{{\color{DarkBlue} y =1.33333 +1.066667 \times x}}

(Round final answer to specified number of decimal places)

Make predictions about happiness values for cookie enthusiasm scores of 3 and 10.

Let x = 3,

y =1.33333 +1.066667 \times x

y =1.33333 +1.066667 \times 3

y =1.33333 +3.2

\hat{y} =4.53333

\mathbf{{\color{DarkOrange} \hat{y} =4.53 }}

(Round final answer to specified number of decimal places)

Let x = 10

y =1.33333 +1.066667 \times x

y =1.33333 +1.066667 \times 10

y =1.33333 +10.66667

y =12.00000

\mathbf{{\color{DarkGreen} \hat{y }=12.00 }}